Q1) You are studying a gene called ‘Muggle’ and the partial sequence of the encoded RNA is given below:

5’ – CGACUAGAUGCAACGUCUUACGAACUAGG-3’

  1. Write down the sequence for the antisense strand of DNA from which this RNA has been transcribed. Make sure to clearly indicate your 5’ and 3’ ends of the sequence.
  2. Can this RNA fragment be used as a probe to detect the location of the ‘Muggle’ gene on the chromosome? Briefly explain your answer• Two organisms have a genome of equal size (equal number of bases in the DNA). The AT content of genome one is 35% and the A
Answer #1

Antisense strand,

3' GCTCATCTACGTTGCAGAATGCTTGATCC 5'

No, the given RNA cannot be used to detect the location of the gene. This is because both gene and RNA are in 5' - 3' direction. They have same sequence, except for T and U. RNA contains U and DNA contains T. Fragments with same sequences do not base pair with each other.

2. Tm = 2 (A+T) + 4 (G+C)

Organism 1 :

AT = 35%

GC = 65%

Tm = 2 (35) + 4 (65) = 330℃

Organism 2 :

AT = 20%

GC = 80%

Tm = 2 (20) + 4 (80) = 360℃

Organism 2 have higher Tm because it has more GC content. Because G and C pair with each other by 3 hydrogen bonds, a higher temperature is required to break this bond as compared to 2 hydrogen bonds of AT.

Please rate high.